# Blade: A Derivative-free Bayesian Inversion Method using Diffusion Priors

### v1

@@ 1 Introduction
Inverse problems, which seek to infer underlying system states or parameters from indirect and noisy observations, are fundamental to numerous applications in science and engineering. For instance, numerical weather prediction requires inferring the atmospheric state by assimilating observational data from weather stations and satellites (Bannister, 2017). Solving these problems requires overcoming three major challenges: first, they are often high-dimensional and ill-posed, meaning that the solution may be non-unique or unstable under perturbations (Hadamard, 2014); second, the design of priors or regularizers is non-trivial and has a significant impact on the solution; third, the associated forward models may involve complicated numerical algorithms that make derivative calculation impractical. Indeed, the weather example admits multiple possible solutions, demands carefully designed priors, and involves intricate numerical procedures with various non-differentiable steps like remapping, branching, and discrete search (Park & Xu, 2013; White, 2000). As such, derivative-free Bayesian inversion methods that use flexible priors to perform reliable uncertainty quantification for high dimensional problems are quite desirable (Park & Xu, 2013).
@@
where ${\bm{y}}\in\mathbb{R}^{m}$ is the observation or measurement, ${\bm{x}}^{*}\in\mathbb{R}^{n}$ is the unknown state, $\mathcal{G}$ is the forward model accessible only via forward evaluations (i.e., black-box access), and $\bm{\epsilon}\in\mathbb{R}^{m}$ is the measurement noise, often modeled as additive Gaussian $\mathcal{N}(0,\sigma_{y}^{2}{\mathbf{I}})$. The Bayesian framework characterizes the solution as a posterior distribution $p({\bm{x}}^{*}|{\bm{y}})\propto p({\bm{x}}^{*})p(y|{\bm{x}}^{*})$ that enables uncertainty quantification for principled decision-making (Sanz-Alonso et al., 2023). As the gradient of $\mathcal{G}$ is difficult or impractical to compute, one typically resorts to derivative-free Bayesian inversion methods.
@@
Traditional methods for derivative-free Bayesian inversion include Markov chain Monte Carlo (MCMC) methods (Geyer, 1992; Gelman et al., 1997; Cotter et al., 2013) and Sequential Monte Carlo (SMC) (Del Moral et al., 2006). These methods offer convergence guarantees but face significant scalability challenges for high dimensional problems. Approximate Bayesian methods (Garbuno-Inigo et al., 2020a; Carrillo et al., 2022; Huang et al., 2022a) offer better efficiency, but often struggle to capture complex posteriors. Additionally, these methods require access to the prior density (up to a normalizing constant), which can be difficult to directly model in high dimensions.
Many recent derivative-free algorithms (Zheng et al., 2025a; Tang et al., 2024; Huang et al., 2024) leverage diffusion models (DMs) as plug-and-play priors for solving high-dimensional inverse problems with complex prior distributions. DMs can flexibly capture complex prior distributions from data, but require optimization or sampling for posterior inference, mainly due to modeling the score function rather than the density. Optimization-based approaches disregard posterior spread and thus often fail to capture spread or uncertainty even in simple Gaussian settings with linear forward models (see Fig. 1). Sampling-based approaches can be asymptotically correct (Trippe et al., 2022; Wu et al., 2023; Cardoso et al., 2024; Dou & Song, 2024), but are typically strictly tailored for linear forward models and do not generalize to the nonlinear setting.
@@ 2 Background
Diffusion Models. We consider diffusion models in the unified EDM framework (Karras et al., 2022). Diffusion models define a forward stochastic process to evolve the original data distribution $p_{0}({\bm{x}})$ to an approximately Gaussian distribution $p_{T}({\bm{x}})=\mathcal{N}(0,s^{2}(T)\sigma(T)^{2}{\mathbf{I}})$, where $\sigma(t)$ is a pre-defined noise schedule function and $s(t)$ is the pre-defined scaling function. Without loss of generality, we set $s(t)=1$ because every other schedule is equivalent to it up to a simple reparameterization as shown in Karras et al. (2022). We consider the following form of denoising diffusion process:
$\mathrm{d}{\bm{x}}_{t}=-\left(2\dot{\sigma}(t)\sigma(t)+\beta(t)\right)\nabla_{{\bm{x}}_{t}}\log p\left({\bm{x}}_{t};\sigma(t)\right)\mathrm{d}t+\left(\sqrt{2\dot{\sigma}(t)\sigma(t)}+\sqrt{2\beta(t)}\right)\mathrm{d}\bar{{\bm{w}}}_{t},$ (2)
where $\beta(t)$ can be any non-negative function as shown in Zhang & Chen (2022). Generating new samples from $p_{0}({\bm{x}})$ amounts to integrating Eq. (2) from a random sample from $p_{T}({\bm{x}}_{T})$. This requires computation of the time-dependent score function $\nabla\log p({\bm{x}}_{t};\sigma(t))$, which can be approximated with a neural network: $s_{\theta}({\bm{x}}_{t},t)\approx\nabla\log p({\bm{x}}_{t};\sigma(t))$. In our work, we assume that we have access to such a pre-trained score function, which we will simply refer to as the diffusion model.
Split Gibbs Sampling. The Split Gibbs Sampler (SGS) (Vono et al., 2019) is a Markov chain Monte Carlo (MCMC) method that aims to sample the posterior $p({\bm{x}}\mid{\bm{y}})\propto p({\bm{y}}\mid{\bm{x}})p({\bm{x}})=\exp(-f({\bm{x}};{\bm{y}})-g({\bm{x}}))$, where $f({\bm{x}};{\bm{y}})=-\log p({\bm{y}}\mid{\bm{x}})$ and $g({\bm{x}})=-\log p({\bm{x}})$. Instead of direct posterior sampling, SGS samples the auxiliary distribution:
@@
where ${\bm{z}}\in\mathbb{R}^{n}$ is an auxiliary variable and $\rho$ is a parameter that controls the distance between $\bm{x}$} and $\bm{z}$}. As shown by Vono et al. (2019), sampling the posterior becomes equivalent to sampling Eq. (3) as $\rho$ approaches 0. Suppose ${\bm{x}}^{(0)}$ is the initial state and $k$ is iteration index. Sampling Eq. (3) is achieved using a Gibbs sampling procedure that alternates between the following two steps:
@@
Diffusion-based Split Gibbs Sampling. Recent works (Bouman & Buzzard, 2023; Coeurdoux et al., 2023; Xu & Chi, 2024; Wu et al., 2024; Chu et al., 2025) have explored adapting generative model priors into the split Gibbs framework. These methods focus on developing different ways to realize the prior step. This is orthogonal to our Blade method, which primarily focuses on the design of the likelihood step. For the prior step, we follows the practice of Wu et al. (2024) which deduces the prior step to a denoising diffusion process within the EDM framework (Karras et al., 2022), enabling the use of any pre-trained diffusion model as the generative prior. For the likelihood step, the prior work either uses gradient-based Langevin Monte Carlo or derives the closed-form expression of $\pi^{Z|X}$ for linear problems, which requires the access to the gradient or adjoint operator of the forward model. In contrast, our Blade method implements the likelihood step as an interacting particle system which does not rely on derivative or adjoint operator information.
Ensemble Kalman Methods. Ensemble Kalman methodology was first introduced by Evensen (1994) as a way of performing statistical linearization (Booton, 1954): estimating a surrogate linear model from samples (e.g., those generated using a black-box forward model). This approach is appealing since it sidesteps the need to compute gradients from the generating process, and can then be integrated into various Bayesian inference schemes. In the inverse problem setting, this idea was used to develop optimization-based frameworks such the Ensemble Kalman Inversion (EKI) framework (Iglesias et al., 2013; Kovachki & Stuart, 2018; Iglesias, 2016; Chada et al., 2020; Huang et al., 2022b) as well as derivative-free diffusion guidance methods (Kim et al., 2024; Zheng et al., 2025a). Our Blade method also builds upon the fundamental idea of statistical linearization of a black-box forward model, and carefully incorporates it into a split Gibbs framework with a diffusion prior to generate well-calibrated posterior samples for high-dimensional inverse problems.
@@ 3.1 Derivative-free likelihood step via statistical linearization
Let ${\mathbf{X}}^{(k)}=\{{\bm{x}}^{(j)}\}_{j=1}^{J}$ denote the ensemble of $J$ particles at $k$-th alternating iteration of the SGS framework. In the likelihood step (Eq. (4)), we aim to sample ${\bm{z}}^{(j)}$ from $\pi^{Z|X={\bm{x}}^{(j)}}({\bm{z}})\propto\exp(-f({\bm{z}};{\bm{y}})-\frac{1}{2\rho^{2}}\|{\bm{z}}-{\bm{x}}^{(j)}\|_{2}^{2})$ for each $j\in\{1,\dots,J\}$. Our starting point is the covariance-preconditioned Langevin dynamics with the large particle limit:
@@
where $C^{-1}_{t}$ is the pseudo-inverse of the covariance matrix. This approach, known as statistical linearization, was first introduced in Booton (1954) and recently used in Kim et al. (2024); Zheng et al. (2025a). Statistical linearization is exact when $\mathcal{G}$ is linear. Let $D\mathcal{G}$ denote the Jacobian of $\mathcal{G}$ and $\tilde{{\bm{z}}}_{t}={\bm{z}}_{t}-\bar{{\bm{z}}}_{t}$. Replacing $D\mathcal{G}$ with $DA_{t}=A_{t}$ in $\nabla f({\bm{z}}_{t}^{(j)};{\bm{y}})$ yields:
@@
$\mathrm{d}{\bm{z}}_{t}^{(j)}=-\left[\frac{1}{\sigma_{y}^{2}}\mathbb{E}_{q_{t}}[\tilde{{\bm{z}}}_{t}(\mathcal{G}({\bm{z}}_{t})-\mathbb{E}_{q_{t}}\mathcal{G}({\bm{z}}_{t}))^{\top}](\mathcal{G}({\bm{z}}_{t}^{(j)})-{\bm{y}})+\frac{1}{\rho^{2}}C_{t}({\bm{z}}_{t}^{(j)}-{\bm{x}}^{(j)})\right]\mathrm{d}t+\sqrt{2C_{t}}\mathrm{d}{\bm{w}}_{t}.$ (9)
Thanks to the covariance preconditioner, the dynamics in Eq. (9) does not require explicit computation of $\hat{A}$ and thus avoids computing the matrix pseudo-inverse. Further, Eq. (9) eliminates the need for derivatives of the forward model, allowing us to run algorithm with only black-box access to $\mathcal{G}$. The dynamics in Eq. (9) shares a similar structure as that of Ensemble Kalman Sampling (EKS) (Garbuno-Inigo et al., 2020a) and ALDI (Garbuno-Inigo et al., 2020b). However, a key distinction is that each particle in Blade has its own target distribution associated with ${\bm{x}}^{(j)}$, enabling multi-modal sampling, whereas EKS and ALDI share the same potential function across all the particles.
@@ 3.2 Ensemble-based prior step via denoising diffusion
Let ${\mathbf{Z}}^{(k)}=\{{\bm{z}}^{(j)}\}_{j=1}^{J}$ denote the ensemble of $J$ particles at $k$-th alternating iteration of the SGS framework. In the prior step (Eq. (5)), we aim to sample ${\bm{x}}^{(j)}$ from $\pi^{X|Z={\bm{z}}^{(j)}}({\bm{x}})\propto\exp(-g({\bm{x}})-\frac{1}{2\rho^{2}}\|{\bm{x}}-{\bm{z}}^{(j)}\|_{2}^{2})$ for each $j\in\{1,\dots,J\}$. As shown in Coeurdoux et al. (2023); Wu et al. (2024), the prior step can be formulated as denoising diffusion process. Specifically, recall that the forward process gives $p({\bm{x}}_{t}\mid{\bm{x}}_{0})=\mathcal{N}({\bm{x}}_{t};{\bm{x}}_{0},\sigma(t)^{2}{\mathbf{I}})$ under the EDM (Karras et al., 2022) framework, where $s(t)=1$. By Bayes’ theorem, we have:
@@ 3.3 Practical algorithm
  • Blade (main) ensures that the invariant measure of the finite-particle system remains the same as that of Eq. (9). As shown in Nüsken & Reich (2019), the covariance-preconditioned stochastic process requires additional correction term as the diffusion term depends on the evolving particle. For $j$-th particle, we add a correction term to the drift of Eq. (9), yielding:
@@
    where $n$ is the dimensionality of $\bm{z}$}, $J$ is the ensemble size. Lemma 3 verifies that Eq. (11) has an invariant measure that is identical to that of Eq. (9). Intuitively, the correction term $\frac{n+1}{J}({\bm{z}}_{t}^{(j)}-\bar{{\bm{z}}}_{t})$ pushes the particles away from each other and vanishes when $J\gg n$. For the computation of $\sqrt{C_{t}}$, we use the construction proposed in Garbuno-Inigo et al. (2020b) where $\sqrt{C_{t}}=\frac{1}{\sqrt{J}}({\bm{z}}_{t}^{(1)}-\bar{{\bm{z}}}_{t},\dots,{\bm{z}}_{t}^{(J)}-\bar{{\bm{z}}}_{t})\in\mathbb{R}^{n\times J}$, which avoids explicit matrix square roots.
@@
Implementation of prior step. Detailed pseudocode for the prior step can be found in Algorithm 3. Note that in Algorithm 3, ${\mathbf{X}}_{i}$ represents the ensemble of particles at $t_{i}$ step, and the updates for all particles can be computed in parallel. For discretization, we use the Euler method with the step size scheme in Karras et al. (2022). Further implementation details are deferred to Appendix B.1.
@@
Like all other ensemble Kalman methods, when the ensemble size is finite, the statistical linearization in a single likelihood step of Blade is confined to a subspace spanned by the particles, meaning that it is typically not a full-rank update, which is sometimes problematic in other ensemble Kalman frameworks. However, since Blade alternates between likelihood and prior steps, the randomness and nonlinearity of the prior step generally implies that sampling trajectory will explore all dimensions (assuming a full-rank prior). To verify this, we empirically track the rank of the ensemble span accumulated over iterations. Fig. 3 shows that the actual rank closely tracks the max possible rank.

### v2

@@ 1 Introduction
Inverse problems, which seek to infer underlying system states or parameters from indirect and noisy observations, are fundamental to numerous applications in science and engineering. For instance, numerical weather prediction requires inferring the atmospheric state by assimilating observational data from weather stations and satellites (bannister2017review). Solving these problems requires overcoming three major challenges: first, they are often high-dimensional and ill-posed, meaning that the solution may be non-unique or unstable under perturbations (hadamard2014lectures); second, the design of priors or regularizers is non-trivial and has a significant impact on the solution; third, the associated forward models may involve complicated numerical algorithms that make derivative calculation impractical. Indeed, the weather example admits multiple possible solutions, demands carefully designed priors, and involves intricate numerical procedures with various non-differentiable steps like remapping, branching, and discrete search (park2013data; white2000ifs). As such, derivative-free Bayesian inversion methods that use flexible priors to perform reliable uncertainty quantification for high dimensional problems are quite desirable (park2013data).
@@
where ${\bm{y}}\in\mathbb{R}^{m}$ is the observation or measurement, ${\bm{x}}^{*}\in\mathbb{R}^{n}$ is the unknown state, $\mathcal{G}$ is the forward model accessible only via forward evaluations (i.e., black-box access), and $\bm{\epsilon}\in\mathbb{R}^{m}$ is the measurement noise, often modeled as additive Gaussian $\mathcal{N}(0,\sigma_{y}^{2}{\mathbf{I}})$. The Bayesian framework characterizes the solution as a posterior distribution $p({\bm{x}}^{*}|{\bm{y}})\propto p({\bm{x}}^{*})p(y|{\bm{x}}^{*})$ that enables uncertainty quantification for principled decision-making (sanz2023inverse). As the gradient of $\mathcal{G}$ is difficult or impractical to compute, one typically resorts to derivative-free Bayesian inversion methods.
@@
Traditional methods for derivative-free Bayesian inversion include Markov chain Monte Carlo (MCMC) methods (geyer1992practical; gelman1997weak; cotter2013mcmc) and Sequential Monte Carlo (SMC) (del2006sequential). These methods offer convergence guarantees but face significant scalability challenges for high dimensional problems. Approximate Bayesian methods (garbuno2020interacting; carrillo2022consensus; huang2022efficient) offer better efficiency, but often struggle to capture complex posteriors. Additionally, these methods require access to the prior density (up to a normalizing constant), which can be difficult to directly model in high dimensions.
Many recent derivative-free algorithms (zheng2024ensemble; tang2024solving; huang2024symbolic) leverage diffusion models (DMs) as plug-and-play priors for solving high-dimensional inverse problems with complex prior distributions. DMs can flexibly capture complex prior distributions from data, but require optimization or sampling for posterior inference, mainly due to modeling the score function rather than the density. Optimization-based approaches disregard posterior spread and thus often fail to capture spread or uncertainty even in simple Gaussian settings with linear forward models (see Fig. 1). Sampling-based derivative-free approaches can be asymptotically correct (trippe2022diffusion; wu2023practical; cardoso2024monte; dou2024diffusion), but have not yet demonstrated the capability to produce probabilistically calibrated samples in the high-dimensional nonlinear setting.
@@ 2 Background
Diffusion Models. We consider diffusion models in the unified EDM framework (karras2022elucidating). Diffusion models define a forward stochastic process to evolve the original data distribution $p_{0}({\bm{x}})$ to an approximately Gaussian distribution $p_{T}({\bm{x}})=\mathcal{N}(0,s^{2}(T)\sigma(T)^{2}{\mathbf{I}})$, where $\sigma(t)$ is a pre-defined noise schedule function and $s(t)$ is the pre-defined scaling function. Without loss of generality, we set $s(t)=1$ because every other schedule is equivalent to it up to a simple reparameterization as shown in karras2022elucidating. We consider the following form of denoising diffusion process:
$\mathrm{d}{\bm{x}}_{t}=-\left(2\dot{\sigma}(t)\sigma(t)+\beta(t)\right)\nabla_{{\bm{x}}_{t}}\log p\left({\bm{x}}_{t};\sigma(t)\right)\mathrm{d}t+\left(\sqrt{2\dot{\sigma}(t)\sigma(t)+2\beta(t)}\right)\mathrm{d}\bar{{\bm{w}}}_{t},$ (2)
where $\beta(t)$ can be any non-negative function as shown in zhang2022fast. Generating new samples from $p_{0}({\bm{x}})$ amounts to integrating Eq. (2) from a random sample from $p_{T}({\bm{x}}_{T})$. This requires computation of the time-dependent score function $\nabla\log p({\bm{x}}_{t};\sigma(t))$, which can be approximated with a neural network: $s_{\theta}({\bm{x}}_{t},t)\approx\nabla\log p({\bm{x}}_{t};\sigma(t))$. In our work, we assume that we have access to such a pre-trained score function, which we will simply refer to as the diffusion model.
Split Gibbs Sampling. The Split Gibbs Sampler (SGS) (vono2019split) is a Markov chain Monte Carlo (MCMC) method that aims to sample the posterior $p({\bm{x}}\mid{\bm{y}})\propto p({\bm{y}}\mid{\bm{x}})p({\bm{x}})=\exp(-f({\bm{x}};{\bm{y}})-g({\bm{x}}))$, where $f({\bm{x}};{\bm{y}})=-\log p({\bm{y}}\mid{\bm{x}})$ and $g({\bm{x}})=-\log p({\bm{x}})$. Instead of direct posterior sampling, SGS samples the auxiliary distribution:
@@
where ${\bm{z}}\in\mathbb{R}^{n}$ is an auxiliary variable and $\rho$ is a parameter that controls the distance between $\bm{x}$} and $\bm{z}$}. As shown by vono2019split, sampling the posterior becomes equivalent to sampling Eq. (3) as $\rho$ approaches 0. Suppose ${\bm{x}}^{(0)}$ is the initial state and $k$ is iteration index. Sampling Eq. (3) is achieved using a Gibbs sampling procedure that alternates between the following two steps:
@@
Diffusion-based Split Gibbs Sampling. Recent works have explored adapting generative model priors into the Gibbs sampling (janati2025mixture; achituveinverse) and split Gibbs framework (bouman2023generative; coeurdoux2023plug; xu2024provably; wu2024principled; chu2025split). These methods focus on developing different ways to realize the prior step. This is orthogonal to our Blade method, which primarily focuses on the design of the likelihood step. For the prior step, we follow the practice of wu2024principled which deduces the prior step to a denoising diffusion process within the EDM framework (karras2022elucidating), enabling the use of any pre-trained diffusion model as the generative prior. For the likelihood step, the prior work either uses gradient-based Langevin Monte Carlo or derives the closed-form expression of $\pi^{Z|X}$ for linear problems, which requires the access to the gradient or adjoint operator of the forward model. In contrast, our Blade method implements the likelihood step as an interacting particle system which does not rely on derivative or adjoint operator information.
Ensemble Kalman Methods. Ensemble Kalman methodology was first introduced by evensen1994sequential as a way of performing statistical linearization (booton1954nonlinear): estimating a surrogate linear model from samples (e.g., those generated using a black-box forward model). This approach is appealing since it sidesteps the need to compute gradients from the generating process, and can then be integrated into various Bayesian inference schemes. In the inverse problem setting, this idea was used to develop optimization-based frameworks such the Ensemble Kalman Inversion (EKI) framework (iglesias2013ensemble; kovachki2018eki; iglesias2016regularizing; chada2020tikhonov; huang2022iterated) as well as derivative-free diffusion guidance methods (kim2024derivative; zheng2024ensemble). Our Blade method also builds upon the fundamental idea of statistical linearization of a black-box forward model, and carefully incorporates it into a split Gibbs framework with a diffusion prior to generate well-calibrated posterior samples for high-dimensional inverse problems.
@@ 3.1 Derivative-free likelihood step via statistical linearization
Let ${\mathbf{X}}^{(k)}=\{{\bm{x}}^{(j)}\}_{j=1}^{J}$ denote the ensemble of $J$ particles at $k$-th alternating iteration of the SGS framework. In the likelihood step (Eq. (4)), we aim to sample ${\bm{z}}^{(j)}$ from $\pi^{Z|X={\bm{x}}^{(j)}}({\bm{z}})\propto\exp(-f({\bm{z}};{\bm{y}})-\frac{1}{2\rho^{2}}\|{\bm{z}}-{\bm{x}}^{(j)}\|_{2}^{2})$ for each $j\in\{1,\dots,J\}$. Our starting point is the covariance-preconditioned Langevin dynamics with the large particle limit, which is known to have improved conditioning and convergence under quadratic potentials (reich2021fokker; garbuno2020interacting):
@@
where $C^{-1}_{t}$ is the pseudo-inverse of the covariance matrix. This approach, known as statistical linearization, was first introduced in booton1954nonlinear and recently used in kim2024derivative; zheng2024ensemble. Statistical linearization is exact when $\mathcal{G}$ is linear. Let $D\mathcal{G}$ denote the Jacobian of $\mathcal{G}$ and $\tilde{{\bm{z}}}_{t}={\bm{z}}_{t}-\bar{{\bm{z}}}_{t}$. Replacing $D\mathcal{G}$ with $DA_{t}=A_{t}$ in $\nabla f({\bm{z}}_{t}^{(j)};{\bm{y}})$ yields:
@@
$\mathrm{d}{\bm{z}}_{t}^{(j)}=-\left[\frac{1}{\sigma_{y}^{2}}\mathbb{E}_{q_{t}}[\tilde{{\bm{z}}}_{t}(\mathcal{G}({\bm{z}}_{t})-\mathbb{E}_{q_{t}}\mathcal{G}({\bm{z}}_{t}))^{\top}](\mathcal{G}({\bm{z}}_{t}^{(j)})-{\bm{y}})+\frac{1}{\rho^{2}}C_{t}({\bm{z}}_{t}^{(j)}-{\bm{x}}^{(j)})\right]\mathrm{d}t+\sqrt{2C_{t}}\mathrm{d}{\bm{w}}_{t},$ (9)
which does not require explicit computation of $\hat{A}$ and thus avoids computing the matrix pseudo-inverse. Eq. (9) also eliminates the need for derivatives of the forward model, allowing us to run the algorithm with only black-box access to $\mathcal{G}$. The dynamics in Eq. (9) shares a similar structure as that of Ensemble Kalman Sampling (EKS) (garbuno2020interacting) and ALDI (garbuno2020affine). However, a key distinction is that each particle in Blade has its own target distribution associated with ${\bm{x}}^{(j)}$, enabling multi-modal sampling, whereas EKS and ALDI share the same potential function across all the particles.
@@ 3.2 Ensemble-based prior step via denoising diffusion
Let ${\mathbf{Z}}^{(k)}=\{{\bm{z}}^{(j)}\}_{j=1}^{J}$ denote the ensemble of $J$ particles at $k$-th alternating iteration of the SGS framework. In the prior step (Eq. (5)), we aim to sample ${\bm{x}}^{(j)}$ from $\pi^{X|Z={\bm{z}}^{(j)}}({\bm{x}})\propto\exp(-g({\bm{x}})-\frac{1}{2\rho^{2}}\|{\bm{x}}-{\bm{z}}^{(j)}\|_{2}^{2})$ for each $j\in\{1,\dots,J\}$. As shown in coeurdoux2023plug; wu2024principled, the prior step can be formulated as denoising diffusion process. Specifically, recall that the forward process gives $p({\bm{x}}_{t}\mid{\bm{x}}_{0})=\mathcal{N}({\bm{x}}_{t};{\bm{x}}_{0},\sigma(t)^{2}{\mathbf{I}})$ under the EDM (karras2022elucidating) framework, where $s(t)=1$. By Bayes’ theorem, we have:
@@ 3.3 Practical algorithm
  • Blade (main) ensures that the invariant measure of the finite-particle system remains the same as that of Eq. (9). As shown in nusken2019note, the covariance-preconditioned stochastic process requires additional correction term as the diffusion term depends on the evolving particle. For $j$-th particle, we add a correction term to the drift of Eq. (9), yielding:
@@
    where $n$ is the dimensionality of $\bm{z}$}, $J$ is the ensemble size, and ${\bm{w}}_{t}\in\mathbb{R}^{J}$. Lemma 3 verifies that Eq. (11) has an invariant measure that is identical to that of Eq. (9). Intuitively, the correction term $\frac{n+1}{J}({\bm{z}}_{t}^{(j)}-\bar{{\bm{z}}}_{t})$ pushes the particles away from each other and vanishes when $J\gg n$. For the computation of $\sqrt{C_{t}}$, we use the construction proposed in garbuno2020affine where $\sqrt{C_{t}}=\frac{1}{\sqrt{J}}({\bm{z}}_{t}^{(1)}-\bar{{\bm{z}}}_{t},\dots,{\bm{z}}_{t}^{(J)}-\bar{{\bm{z}}}_{t})\in\mathbb{R}^{n\times J}$, which avoids explicit matrix square roots.
@@
Implementation of prior step. Detailed pseudocode for the prior step can be found in Algorithm 3. Note that in Algorithm 3, ${\mathbf{X}}_{i}$ represents the ensemble of particles at $t_{i}$ step, and the updates for all particles can be computed in parallel. For discretization, we use the Euler method with the step size scheme in karras2022elucidating. Further implementation details are deferred to Appendix B.1.
@@
Like all other ensemble Kalman methods, when the ensemble size is finite, the statistical linearization in a single likelihood step of Blade is confined to a subspace spanned by the particles, meaning that it is typically not a full-rank update, which is sometimes problematic in other ensemble Kalman frameworks. However, since Blade alternates between likelihood and prior steps, the randomness and nonlinearity of the prior step generally implies that sampling trajectory will explore all dimensions (assuming a full-rank prior). To empirically assess whether Blade suffers from the same limitation, we track the dimensionality of explored space over iterations in the setting where the ensemble size (512) is significantly smaller than the ambient dimension (16384). At iteration $k$, we form the matrix whose columns are the concatenated particles from all iterations up to $k$: $X^{0},Z^{0},\dots,X^{(k)},Z^{(k)}]$. The accumulated rank is the rank of this matrix. Fig. 3 shows that the actual rank closely tracks the max possible rank and it reaches the ambient dimension within 35 steps, indicating that Blade has sufficient posterior exploration without the low-dimensional confinement that limits other ensemble methods.